# Moving Object Segmentation: All You Need Is SAM (and Flow)

### v1

@@ 8 Conclusion
Acknowledgements
This research is supported by the UK EPSRC CDT in AIMS (EP/S024050/1), the Clarendon Scholarship, and the UK EPSRC Programme Grant Visual AI (EP/T028572/1).
@@ Acknowledgements
This Appendix consists of the following sections: (i) In Appendix 0.A, we provide implementation details regarding architectural designs and experimental settings; (ii) In Appendix 0.B, we conduct comprehensive ablation studies for both frame-level and sequence-level segmentation; (iii) In Appendix 0.C (combining motion segmentation with motion-guided segmentation), we discuss the rationale of combining results from FlowI-SAM and FlowP-SAM into a single prediction with some example cases; (iv) In Appendix 0.D, additional quantitative comparisons are provided; (v) In Appendix 0.E, visualisations and failure cases are discussed.
@@ 0.A.2 Sequence-Level Association
In this section, we present the code snip used for sequence-level association (Algorithm 2). We will release the full code upon publication.
@@ 0.B.2 Frame-Level Segmentation: FlowP-SAM
Table 8: Quantitative comparison when the centroid of each groundtruth object mask is provided as a point prompt. Both SAM and FlowP-SAM are evaluated following this setting. The results are shown for frame-level predictions, and our default FlowP-SAM utilises the $1^{\text{st}}$ mask token.
Comparison with FlowI-SAM + SAM [12]. An alternative method for moving object segmentation is to apply SAM to refine the flow-predicted masks by FlowI-SAM. However, such a two-stage method prevents the interaction between RGB and flow information, leading to inferior performance compared to the end-to-end FlowP-SAM, as can be seen in Table 9.
@@
Table 9: Comparison with a two-stage method. The results are shown for frame-level predictions.
@@ 0.B.3 Sequence-Level Mask Association
We observe that (i) Hungarian matching alone presents a strong baseline for RGB-based tracking (using FlowP-SAM + FlowI-SAM), where frame-level predictions are largely consistent; (ii) However, in flow-only cases (using FlowI-SAM) where object identities might disappear, Hungarian matching is prone to lose track of the object. This issue of losing track can be solved using our temporal consistency method, where choosing to propagate appropriate masks helps maintain object permanence. We observe that flow-only method (using FlowI-SAM) benefits considerably from this. We postulate further that using more frames may help even further, but this will involve computing new optical flows, whereas we only currently use those that are already used as input to FlowI-SAM and/or FlowP-SAM.
@@ Appendix 0.D Quantitative Results
FlowP-SAM+FlowI-SAM (seq) $\checkmark$ $\checkmark$ $\mathbf{75.2}$ $\mathbf{74.7}$ $\mathbf{75.7}$ $\mathbf{75.0}$ $\mathbf{74.3}$ $\mathbf{75.6}$ $\mathbf{73.6}$ $\mathbf{71.0}$ $\mathbf{76.1}$

### v2

@@ 8 Discussion
8.0.1 Acknowledgments.
[…]
@@
This Appendix consists of the following sections:
(i) In Appendix 0.A, we provide implementation details regarding architectural designs and experimental settings;
(ii) In Appendix 0.B, we conduct comprehensive ablation studies for both frame-level and sequence-level segmentation;
(iii) In Appendix 0.C (combining motion segmentation with motion-guided segmentation), we discuss the rationale of combining results from FlowI-SAM and FlowP-SAM into a single prediction with some example cases;
(iv) In Appendix 0.D, additional quantitative comparisons are provided, along with a discussion on evaluating unsupervised object discovery using DAVIS17.
(v) In Appendix 0.E, visualisations and failure cases are discussed.
@@ 0.A.1 Frame-Level Segmentation
Inference time analysis. As mentioned in Sec. 8 in the main text, the heavy image encoder in SAM leads to extended running times. We conduct a detailed inference time analysis and list the results of our default FlowI-SAM and FlowP-SAM in the first row of Table 5. We also report the potential speed-ups of our method when adapting more efficient image encoders from more recent SAM-like models.
FlowI-SAM FlowP-SAM
Encoder (Flow) Running time s/img Encoder (Flow $\mid$ RGB) Running time s/img
SAM-B $0.46$ SAM-B $\mid$ SAM-H $0.97$
EfficientSAM-S [49] $0.07$ EfficientSAM-S [49] (both) $0.08$
SAM2-B+ [36] $0.08$ SAM2-B+ [36] $\mid$ SAM2-L [36] $0.16$
Table 5: Inference time analysis. Our models can benefit from more efficient image encoders, with the potential speed-ups shown in last two rows.
@@ 0.A.2 Sequence-Level Association
In this section, we present the code snip used for sequence-level association (Algorithm 2).
@@ 0.B.1 Frame-Level Segmentation: FlowI-SAM
$1,$$1$ $64.5$ $78.0$
$2,$$2$ $65.3$ $77.7$
$1,$-$1,2,$-$2$ $65.7$ $79.1$
Table 6: The frame gaps of input flows to FlowI-SAM. The SAM ViT-H image encoder is adopted. The results are shown for frame-level predictions.
Taking max $65.0$ $78.2$
Averaging $65.7$ $79.1$
@@
Figure 6: Qualitative comparison between IoUs estimated by SAM and scores predicted by FlowP-SAM for selecting masks of foreground objects. Only masks with top-5 scores are shown. The masks with higher scores are brighter and arranged to the front.
@@ 0.B.2 Frame-Level Segmentation: FlowP-SAM
Table 9: Quantitative comparison between SAM and FlowP-SAM. The centroid of each groundtruth object mask is provided as a point prompt. The results are shown for frame-level predictions, and our default FlowP-SAM utilises the $1^{\text{st}}$ mask token.
Furthermore, our proposed scores (MOS and fIoU) serve as more accurate indicators of complete foreground objects, compared to the original IoU scores estimated by SAM. Fig. 6 illustrates a qualitative example, where simply ranking the masks with higher IoU scores would result in the selection of background masks (e.g., the ground) and incomplete masks (e.g., the middle pig with an ear missing). In contrast, the averaged MOS + fIoU scores help to correctly identify the complete foreground object masks, which verifies our claim that the proposed scores capture the “objectness” of masks.
Comparison with SAM2 [36]. Following the comparison setup in Fig. 6 (i.e., prompting by a single ground truth centroid), we compare our FlowP-SAM with more recent SAM2. As shown in Table 10, though SAM2 gets noticeably improved from SAM, our FlowP-SAM still demonstrates superior object segmentation performance.
Model DAVIS17-m $\mathcal{J}\uparrow$    DAVIS17 $\mathcal{J}\uparrow$    DAVIS16 $\mathcal{J}\uparrow$
SAM [13] $71.4$ $68.0$ $73.6$
SAM2 [36] $79.4$ $69.7$ $84.0$
 FlowP-SAM $80.0$ $73.4$ $85.6$
Table 10: Quantitative comparison across SAM, SAM2, and FlowP-SAM. The centroid of each groundtruth object mask is provided as a point prompt. The best results among the four output channels are shown.
Comparison with flow-based method + SAM [13]. An alternative method for moving object segmentation is to apply SAM to refine the flow-predicted masks (e.g., by OCLR-flow and by FlowI-SAM). However, such a two-stage method prevents the interaction between RGB and flow information, leading to inferior performance compared to the end-to-end FlowP-SAM, as can be seen in Table 11.
@@
Table 11: Comparison with two-stage methods. The results are shown for frame-level predictions.
@@
Optical Flow Estimation Method. Different from FlowI-SAM results, Table 13 demonstrates a lower impact of flow quality on FlowP-SAM. This is because in FlowP-SAM, the optical flow is mainly adopted to prompt the region of moving objects, while its boundary accuracy only affects the final results marginally.
MaskFlownet [60] $65.3$ $78.9$
RAFT [41] $65.7$ $79.1$
[…]
@@ 0.B.3 Sequence-Level Mask Association
We observe that (i) Hungarian matching alone presents a strong baseline for RGB-based tracking (using FlowP-SAM + FlowI-SAM), where frame-level predictions are largely consistent; (ii) However, in flow-only cases (using FlowI-SAM) where object identities might disappear, Hungarian matching is prone to lose track of the object. This issue of losing track can be solved using our temporal consistency method, where choosing to propagate appropriate masks helps maintain object permanence. We observe that flow-only method (FlowI-SAM) benefits considerably from this. We postulate further that using more frames may help even further, but this will involve computing new optical flows, whereas we only currently use those already adopted as input to FlowI-SAM and/or FlowP-SAM.
@@ Appendix 0.D Quantitative Results
UVOSAM [59] $\checkmark$ -- -- -- -- -- -- $\mathbf{80.9}$ $\mathbf{77.5}$ $\mathbf{84.9}$
FlowP-SAM+FlowI-SAM (seq) $\checkmark$ $\checkmark$ $\mathbf{75.2}$ $\mathbf{74.7}$ $\mathbf{75.7}$ $\mathbf{75.0}$ $\mathbf{74.3}$ $\mathbf{75.6}$ $73.6$ $71.0$ $76.1$
@@
Discussion on DAVIS17 annotations. It is important to note that we found the DAVIS17 dataset may not be suitable for unsupervised VOS evaluations due to issues with object definitions. This can be understood in two aspects: (i) Ambiguous object definitions within the dataset due to class-based annotations. For example, a helmet on a cyclist is not considered a separate object, whereas a smartphone or handbag held by a person is. Such annotations pose a problem for our class-agnostic object discovery method; (ii) Inconsistencies between annotations and motion information, where objects moving together are labeled separately based on their classes. This contradiction can confuse our model, as we use motion information as prompts for object discovery. This is also the underlying reason why we adopt DAVIS17-m (and YTVOS18-m) for evaluation, as they provide a clearer definition of objectness based on (joint) movements.